# Unsupervised Elicitation of Language Models

### v1

@@ 7 Discussion
The role of logical consistency. At first glance, ICM might look like a consistency-based algorithm, and consistency is indeed part of our scoring function (Sec. 2.2). However, as Sec. 5 shows, removing consistency in our scoring function often does not degrade the maximal performance, but increases the variance. Specifically, the algorithm becomes more likely to collapse into degenerate solutions (that have low logical consistency), like assigning the same label to all data points. Therefore, we understand mutual predictability as the most important term that leads to our empirical success. In particular, mutual predictability also likely enforces complex (probabilistic) consistencies, which cannot be easily captured by general axiomatic logical checks.
Unsupervised elicitation as an alignment method. In practice, when using unsupervised elicitation for alignment, we would still need humans in the loop for various parts of the post-training process. For example, ICM can be directly applied to enhance constitutional AI [5] for aligning LMs. Specifically, for each human-specified constitution, we can replicate our pipeline in Sec. 4.4: use ICM to label which assistant response follows the constitution more accurately and train an unsupervised reward model, then use reinforcement learning to optimize and align the assistant towards the constitution. Additionally, we still need humans to validate whether the model is interpreting the constitution as intended, for example using scalable oversight techniques [26, 22, 30].
Limitations. Our algorithm has two important limitations: (1) As shown in Sec. 4.2, it cannot elicit any concepts or skills unless they are “salient” to the pretrained model. (2) It doesn’t work with long inputs because we need to fit many dataset examples into the model’s effective context window when calculating the scoring function, particularly for the mutual predictability term.
@@
Acknowledgments
We would like to thank Alec Radford, Akbir Khan, Monte MacDiarmid, Fabien Roger, John Schulman, Lijie Chen, Ruiqi Zhong, and Jessy Lin for their valuable feedback.
@@ Acknowledgments
Appendix A Additional Implementation Details
A.1 Hyperparameters
@@
A.2 Data Statistics
Table 2 shows the size of train/test splits used for the experiments in Sec. .
@@ A.2 Data Statistics
Appendix B Compute Costs
ICM is one form of inference-time scaling. We thus investigate how many forward passes we need to label each datapoint on average. Specifically, we report the statistics based on labeling $n=128$ datapoints. As shown in Table 3, ICM often requires 2 to 3 forward passes to label each datapoint.
Table 3: The average number of forward passes required to label each datapoint with ICM.
Dataset Avg. # Forward
@@ Appendix B Compute Costs
Appendix C Human Annotation

### v2

@@ 6 Discussion
The role of logical consistency. As Sec. 5 shows, removing consistency often does not degrade the maximal performance, but increases the variance. Specifically, the algorithm becomes more likely to collapse into degenerate solutions that have low logical consistency, like assigning the same label to all data points. Therefore, we understand mutual predictability as the most important term that leads to our empirical success. In particular, mutual predictability also likely enforces complex (probabilistic) consistencies, which cannot be easily captured by general axiomatic logical checks.
Unsupervised elicitation as an alignment method. In practice, when using unsupervised elicitation for alignment, we would still need humans in the loop for various parts of the post-training process. For example, ICM can be directly applied to enhance constitutional AI (Bai et al., 2022b) for aligning LMs. Specifically, for each human-specified constitution, we can replicate our pipeline in Sec. 4.4: use ICM to label which assistant response follows the constitution more accurately and train an unsupervised reward model, then use reinforcement learning to optimize and align the assistant towards the constitution. Additionally, we still need humans to validate whether the model is interpreting the constitution as intended, for example using scalable oversight techniques (Saunders et al., 2022; McAleese et al., 2024; Wen et al., 2024a).
Limitations. As shown in Sec. 4.2, our algorithm cannot elicit any concepts or skills unless they are “salient” to the pretrained model. In addition, one potential concern is that unsupervised elicitation might be related to data contamination from pretraining. While we cannot directly verify this concern as Llama pre-training corpus is not accessible, there are several pieces of evidence that make data contamination less worrying. For example, the production assistant training data in Sec. 4.4 is certainly not involved in Claude 4 Sonnet’s pretraining corpus. See Appendix J for more discussion.
@@
Ethics Statement
While this paper proposes an unsupervised algorithm to elicit superhuman capabilities from LMs, this does not necessarily mean humans will lose control over LMs. As discusssed in Sec. 6 and empirically showed in Sec. 4.4, our method could be combined with human-specified constituions to potentially align powerful LMs with human values.
@@
Appendix A The Use of LLMs
We only use LLMs to polish paper writing. We did not use LLMs to generate experimental code or directly generate the paper draft.
Appendix B Related Work
Scaling Beyond Human Supervision. Recent work has shown diverse failure modes of post-training with unreliable human supervision. For example, LMs can learn to reward-hack human-designed supervision signals (Baker et al., 2025) or even humans themselves (Wen et al., 2024b). To scale beyond human supervision, one standard method is to use high-quality verifiable rewards. For example, in math, we can match model outputs with existing golden solutions (Guo et al., 2025). Unfortunately, such verifiable rewards are unavailable for most tasks. In contrast, our method can provide superhuman-level supervision in broad tasks, even including creating a general helpful, harmless, and honest assistant.
Evidence of Latent Capabilities in LMs. Recent work shows that pre-trained base models have already learned strong capabilities for downstream tasks, and post-training in fact does not add much. For example, pretrained models can achieve a comparable or even higher pass@$k$ than their post-trained counterparts when $k$ is large enough, even when post-training is done with verifiable rewards (Yue et al., 2025). Similarly, pretrained and post-trained models perform nearly identically in decoding, while most distribution shifts occur with stylistic tokens such as discourse markers (Lin et al., 2023). When inspecting model latent representations, recent work also finds that LMs encode strong signals of reasoning correctness (Zhang et al., 2025) or hallucination (Kadavath et al., 2022; Ferrando et al., 2024). Our work dmeonstrates that it is promising to elicit these LMs’ important latent capabilities in a fully unsupervised way.
Unsupervised Elicitation of LMs. CCS (Burns et al., 2022) is one of the most representative works for unsupervised elicitation, which works by solely using simple logical consistency to find latent knowledge. While moderately outperforming the zero-shot prompting baseline, CCS still significantly underperforms supervised approaches. As argued in (Farquhar et al., 2023), CCS, as well as other unsupervised approaches, often cannot find knowledge, because there are many other prominent features that can satisfy logical consistency properties. Our method addresses this challenge by introducing mutual predictability.
Several concurrent studies explore unsupervised elicitation by minimizing label entropy (Zhao et al., 2025; Agarwal et al., 2025), differing from our scoring function. Empirically, these studies focus on math or coding domains using specific Qwen pretrained models. In contrast, our work demonstrates for the first time that unsupervised elicitation algorithms can match or exceed human supervision across pretrained models and a variety of crisp and fuzzy tasks — even including training a general-purpose assistant.
Unsupervised elicitation can also be thought of as a special case of weak-to-strong generalization (Burns et al., 2023; Hase et al., 2024): while they try to use weak human supervision to elicit strong LMs, we seek to ignore the weak human supervision altogether.
Appendix C Illustrative Example of Scoring Function
A higher score indicates that the labels are more logically consistent and mutually predictable according to the pre-trained base model.
To provide more intuition about how our scoring function works, consider a small dataset D consisting of four examples:
  1. $x_{1}$: lying is more harmless than honesty
  2. $x_{2}$: honesty is more harmless than lying
  3. $x_{3}$: hacking unit tests is more harmless than writing correct code
  4. $x_{4}$: writing correct code is more harmless than hacking unit tests
Logical consistency constraints that $A>B$ and $B>A$ cannot both be True, so $y_{1}\neq y_{2}$, $y_{3}\neq y_{4}$. According to the model’s understanding of the harmlessness concept, $x_{2}$ and $x_{4}$ are mutually predictable, and $y_{2}=y_{4}=True$ is more likely than $y_{2}=y_{4}=False$. So the label assignment $\{False,True,False,True\}$ achieves high mutual predictability and logical consistency, yielding a high overall score.
Appendix D Additional Implementation Details
We use Llama 3.1 8B, Llama 3.1 70B, and Claude 4 Sonnet in our experiments. Unless stated otherwise, we always use pretrained models that have received no additional training, i.e. no supervised fine-tuning on demonstrations, RLHF, RL on outcomes, or any other post-training.
D.1 Hyperparameters
@@
For many-shot prompting, we use as many examples as possible that can fit into the model’s context, e.g., 160 examples for Alpaca. For fine-tuning, we train the model for 3 epochs. Specifically, for Llama 8B, we do full parameter fine-tuning with a learning rate of 1e-5; for Llama 70B, we do LoRA fine-tuning with a rank of 16 and a learning rate of 5e-5.
D.2 Data Statistics
Table 2 shows the size of train/test splits used for the experiments in Sec. 4.1.
@@ D.2 Data Statistics
Appendix E Additional Baselines
In this section, we compare ICM to several additional baselines.
E.1 Distillation
We use zero-shot prompting with commericial LMs to generate labels and train models. Note that this baseline has unfair advantages in paramter size and access to external supervision (since these commercial LMs are heavily post-trained on human labels).
Specifically, following prior work (Huang et al., 2022; Prasad et al., 2024; Jiao et al., 2024), for each example, we use GPT-4o to sample $K=10$ labels and do majority-voting to decide the final label. We then fine-tune Llama models on these labels. We show the results in Table 3.
On all benchmarks, fine-tuning on GPT-4o generated labels underperforms our unsupervised algorithm. In particular, on Alpaca, it achieves similar performance to fine-tuning on real human labels, potentially suggesting that commericial post-trained models’ capability in judging helpfulness and harmlessness is bottlenecked by its post-training human data.
Table 3: Our unsupervised algorithm that is solely based on Llama models outperforms model distillation from GPT-4o.
Benchmark Method Accuracy
GSM8K Golden Label $78.1\pm 0.5$
GPT-4o generated label $75.1\pm 0.7$
Ours $77.0\pm 0.8$
TruthfulQA Golden Label $92.0\pm 1.0$
GPT-4o generated label $81.9\pm 1.6$
Ours $90.9\pm 0.6$
Alpaca Human Label $65.5\pm 0.6$
GPT-4o generated label $65.2\pm 0.5$
Ours $68.0\pm 0.7$
Gender Prediction Golden Label $80.5\pm 0.3$
GPT-4o generated label $77.0\pm 0.0$
Ours $79.7\pm 0.4$
E.2 CCS
For each benchmark, we train a linear probe on model activations using CCS (Burns et al., 2022) with the same hyperparameters as in the original paper. Because the CCS loss function does not specify which probe direction corresponds to true or false, we report the maximum accuracy between the two possible directions for each dataset, as in (Burns et al., 2022).
As shown in Table 4, on three benchmarks, ICM outperforms CCS by a large margin.
Table 4: Our unsupervised algorithm outperforms CCS by a large margin. For each benchmark, we report the maximum CCS probe accuracy across layers and between the two possible probe directions.
Benchmark Method Accuracy
GSM8K CCS $67.0\pm 0.001$
Ours $77.0\pm 0.8$
TruthfulQA CCS $63.0\pm 0.001$
Ours $90.9\pm 0.6$
Alpaca CCS $53.0\pm 0.003$
Ours $68.0\pm 0.7$
The performance of CCS is also sensitive to the layer from which activations are taken. We show the benchmark performance for different layers in Figure11.
Figure 11: CCS probe performance varies significantly by layer. We report the maximum CCS probe accuracy between the two possible probe directions.
Appendix F Compute Costs
ICM is one form of inference-time scaling. We thus investigate how many iterations we need to label each datapoint on average. Specifically, we report the statistics based on labeling $n=128$ datapoints. As shown in Table 5, ICM often requires 2 to 3 iterations to label each datapoint.
Table 5: The average number of iterations required to label each datapoint with ICM.
Dataset Avg. # Iteration
@@ Appendix F Compute Costs
Appendix G Scalable ICM
Our algorithm 1 has two scalability limitations. First, it measures mutual predictability with in-context learning and thus requires all labeled examples to fit in the model’s context window. However, for production assitant training data, each example could take thousands of tokens. Second, it sequentially labels one example at a time, which is inefficient. To overcome these limitations, we propose a scalable variant of ICM (Algorithm 3): it uses fine-tuning to measure mutual predictability, and labels examples in parallel batches.
Algorithm 3 Scalable Internal Coherence Maximization
1:Unlabeled Dataset $D_{\text{unlabel}}=\{x_{i}\}$. Labeled Dataset $D=\emptyset$. Pretrained model $\theta$. Number of folds $F$. Number of iterations $G$.
2:Labeled Dataset $\{x_{i},y_{i}\}$.
3:Label $D_{\text{unlabel}}$ with $\theta$: $D\leftarrow\theta(D_{\text{unlabel}})$. $\triangleright$ Initialize
[…]
5:for $g=1,\cdots,G$ do
6:  Partition $D$ randomly into $F$ disjoint folds $\{D_{f}\}$ such that consistency groups remain in the same fold.
7:  for $f=1,\cdots,F$ do
8:   $\hat{\theta}_{f}\leftarrow\text{Train}(\theta,D\setminus D_{f})$.
9:   Relabel $\hat{D}_{f}=\hat{\theta}_{f}(D_{f})$. $\triangleright$ Increase mutual predictability
10:   $\hat{D}_{f}=\texttt{consistencyfix\_maxprob}(\hat{D}_{f})$. $\triangleright$ Resolve relabeling inconsistency
11:  end for
12:  Merge new labels from different folds: $D\leftarrow\bigcup_{f}\hat{D}_{f}$. $\triangleright$ Update labels
13:  Train $\theta$ on updated labels: $\theta\leftarrow\text{Train}(\theta,D)$.
14:end for
Algorithm 4 ConsistencyFix-MaxProb
1:Labeled Dataset $D$. Pretrained model $\theta$. Consistency groups $\{C_{j}\}$, which is a partition of $D$.
2:Updated Labeled Dataset $D$.
3:for $j$ do
4:  $(x^{*},y^{*})=\arg\max_{(x_{i},y_{i})\in C_{j}}P_{\theta}(y_{i}\mid x_{i})$ $\triangleright$ Most confident prediction
5:  for $(x_{i},y_{i})\in C_{j}$ do
6:   $\hat{y}_{i}=\arg\max_{y}c(x_{i},y,x^{*},y^{*})$ $\triangleright$ Enforce consistency
7:   $D\leftarrow D\cup\{(x_{i},\hat{y}_{i})\}$
8:  end for
9:end for
Measure mutual predictability. First, to overcome the context window limitation, we replace in-context learning with fine-tuning. However, since mutual predictability is based on the probability of each label conditioned on all other |D|1|D|-1 labels, measuring it directly would require fine-tuning $|D|$ individual models, which is expensive as we scale up $D$. To improve efficiency, we approximate conditioning on all but one label with conditioning on all but a few labels. This allows multiple labels to share the same set of conditioned examples, and thus the same fine-tuned model when measuring mutual predictabilility.
Specifically, we randomly partition $D$ into $F$ disjoint subsets, i.e., $D=\cup D_{1},\cdots D_{F}$. Let $t_{i}$ denote the subset that $(x_{i},y_{i})$ belongs to. We approximate $P_{\theta}(y_{i}|x_{i},D\setminus(x_{i},y_{i}))$ with $P_{\theta}(y_{i}|x_{i},D\setminus D_{t_{i}})$.
To search for mutually predictable labels, for each fold $D_{f}$, we train one model on $D\setminus D_{f}$ and use it to relabel examples in $D_{f}$. In this way, searching for mutually predictable labels only needs $|D|/F$ finetuning runs (parallel) and $|D|$ parallel zero-shot inference.
Enforce logical consistency. Algorithm 1 fixes inconsistency by assigning the consistent labeling that achieves highest scores, i.e., maximizing mutual predictability and consistency. However, measuring the mutual predictability for every consistent labeling is expensive: it requires separate fine-tuning on each consistent labeling. We introduce a simpler algorithm to fix inconsistency. For each consistency group, it first identifies the examples where the model’s prediction is most confident, and then adjusts the labels on other examples in the same consistency group to be consistent with it (Algorithm 4).
Appendix H Evaluation Results of Claude 4 Sonnet Reward Models
We experiment with Claude 3.5 Haiku and Claude 4 Sonnet to study how ICM scales with model size, and experiment with training on 512 and 5K preference pairs to study how ICM scales with unlabeled data size. As baselines, we train human-supervised RMs with the same models on the same data but with production-grade human labels.
We evaluate reward models (RMs) on Rewardbench (Lambert et al., 2024). Results are shown in Figure 12. ICM scales well with model sizes: the average performance on RewardBench increases from 0.63 to 0.74 when training on 512 examples, from 0.69 to 0.79 when training on 5K examples.
Comparing our unsupervised algorithm with human supervision, we have the following findings.
In a low-data regime where human labels are too expensive to collect [^3] We are particularly interested in eliciting capabilites on these challenging tasks in this paper, as using AIs to assist humans on these tasks would be highly valuable., ICM outperforms human supervision by a large margin. For example, our unsupervised RM trained on 5K unlabeled data outperforms training on 512 human labels by 15.2% with Claude 3.5 Haiku, and by 6% with Claude 4 Sonnet on average. Note that unsupervised algorithm can be trained on unlimited amount (>>5K>>5K) of unlabeled data, so the performance gain is likely to further improve.
If collecting thousands of human labels is plausible, results depend on model capabilities. For weak models like Claude 3.5 Haiku, ICM can slightly outperform training with human supervision. However, for strong models like Claude 4 Sonnet, ICM underpeforms training with human supervision on average. Taking a closer look at the comparison results across each test set, we find that on two challenging test sets (Chat-hard and Math-debiased), while most RMs achieve near-random accuracy, the Claude 4 Sonnet-based RM trained on 5K human labels achieve a substantially higher accuracy of 0.66. Overall, since the performance of unsupervised algorithms would be bottlenecked by LMs’ existing latent capababilities, it is unsurprising that unsupervised algorithms would underpeform training on high-quality external labels in certain cases (e.g. on crisp tasks like mathematical reasoning). However. for future LMs that have broad superhuman capabilities on our interested tasks, we still expect unsupervised algorithms to beat human supervision baselines.
Figure 12: Evaluating the accuracy of reward models on RewardBench. Unsupervised RM is trained with our algorithm, while the human-supervised RM is trained with production-grade human labels.
Appendix I Human Annotation
@@
Appendix J Discussion: Data Contamination
While we cannot directly check data contamination since we don’t have access to Llama pre-training corpus, there are several pieces of evidence that make data contamination less worrying.
  1. 1.
    As shown in Figure 3, the zero-shot performance of Llama base models are close to randomly guessing (e.g. 60% on TruthfulQA, 50% on ALpaca, and 48% on GSM8K)
  2. 2.
    We reformat GSM8K and TruthfulQA into classification tasks, which is differnt from the original data.
  3. 3.
    Most of our experiments are based on llama models. show that while Qwen models have serious data leakage issue that make even optimizing with random rewards increases their performance on benchmarks, Llama models do not.
  4. 4.
    In Sec. 4.4, the production assistant training data is not involved in the pre-training corpus of Claude 4 Sonnet.